OpenStudy (anonymous):

Find a power series that have an interval of convergence of [1,3].

OpenStudy (misty1212):

HI!!

OpenStudy (anonymous):

hello :D

OpenStudy (misty1212):

probably easiest just to adjust \(\frac{1}{1-x}\) which has radius of convergence \(|x<1\)

OpenStudy (anonymous):

move \[\sum\frac{x^n}{n}\] to the right one unit

OpenStudy (anonymous):

oh no that is wrong sorry

OpenStudy (anonymous):

move it to the right 2 units

OpenStudy (anonymous):

Would it be series of (x+2)^n / n then?

OpenStudy (anonymous):

no that moves it to the LEFT (don;t forget pre calc!)

OpenStudy (anonymous):

Oh oops yeah, so it would be (x-2)^n / n!

OpenStudy (anonymous):

yeah that would work get \(|x-2|<1\) solve and get what you want